How many ounces of pure bran flakes must be added to 60 oz of cereal that is 20% bran flakes to produce a mixture that is 40% bran flakes?

Found 2 solutions by stanbon, ewatrrr:
Answer by stanbon(75887)

How many ounces of pure bran flakes must be added to 60 oz of cereal that is 20% bran flakes to produce a mixture that is 40% bran flakes?
------------------
Equation:
bran + bran = bran
0.20*60 + 1.00x = 0.40(60+x)
---
Multiply thru by 100 to get:
20*60 + 100x = 40*60 + 40x
60x = 20*60
x = 20 ounces (amt. of pure bran needed in the mixture)
--------------------
Cheers,
Stan H.

Answer by ewatrrr(24785)


Hi,

Let x represent the amount of pure bran flakes (an 100% solution)
.20*60oz + 1.00x = .40(60oz + x)
solving for x
.60x = .20*60oz
x = 20oz of oure bran flakes need to be added
checking our answer
12 + 20 = .40*80 = 32